Solve the following inequality. Write the answer in interval notation. If the answer includes more than one interval write the intervals separated by the "union" symbol, U. If needed enter as "infinity".
1[3x+5]+5<=30
[ and ] are verical bars :)
Please help me!

Answer by jim_thompson5910(35256)

1|3x+5|+5 <= 30

1|3x+5| <= 30-5

1|3x+5| <= 25

|3x+5| <= 25/1

|3x+5| <= 25

3x+5 <= 25 and 3x+5 >= -25

-25 <= 3x+5 <= 25

-25-5 <= 3x <= 25-5

-30 <= 3x <= 20

-30/3 <= x <= 20/3

-10 <= x <= 20/3

So the answer in interval notation is

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------